what is the exterior angle and number of sides if the interior equal 140 degrees
Answer by Alan3354(69443)

what is the exterior angle and number of sides if the interior equal 140 degrees
-------
Ext angle = 180 - 140 = 40 degs
n = 360/40 = 9 sides